Find an equation of the line containing the given pair of points. (-7, -8) and (-3, -1)
Answer by mananth(16949)

x1 y1 x2 y2
-7 -8 -3 -1

slope m =(y2-y1)/(x2-x1)
( -1 - -8 )/( -3 - -7 )
( 7 / 4 )
m= 7/4
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( -7 , -8 ) in
Y = m x + b
-8.00 = 7/4 * -7 + b
-8.00 = -49/4 + b
b= -8 + 17/4
b= -12 1/ 4
So the equation will be
Y = 7/4 x + 17/4
m.ananth@hotmail.ca